# Overcoming the Stability Gap in Continual Learning

### v3

@@ Abstract
Pre-trained deep neural networks (DNNs) are being widely deployed by industry for making business decisions and to serve users; however, a major problem is model decay, where the DNN’s predictions become more erroneous over time, resulting in revenue loss or unhappy users. To mitigate model decay, DNNs are retrained from scratch using old and new data. This is computationally expensive, so retraining happens only once performance has significantly decreased. Here, we study how continual learning (CL) could potentially overcome model decay in large pre-trained DNNs and also greatly reduce computational costs for keeping DNNs up-to-date. We identify the “stability gap” as a major obstacle in our setting. The stability gap refers to a phenomenon where learning new data causes large drops in performance for past tasks before CL mitigation methods eventually compensate for this drop. We test two hypotheses for why the stability gap occurs and identify a method that vastly reduces this gap. In large-scale experiments for both easy and hard CL distributions (e.g., class incremental learning), we demonstrate that our method reduces the stability gap and greatly increases computational efficiency. Our work aligns CL with the goals of the production setting, where CL is needed for many applications.
@@ 1 Introduction
Continual learning (CL) is a promising solution for preventing model decay (Huyen, 2022a; b; Jain & Shenoy, 2022), where in CL the goal is to update a DNN incrementally with new data while preserving the previous knowledge (Parisi et al., 2019). In a GrubHub study, this enabled them to avoid model decay and provided a 45$\times$ decrease in training costs compared to daily offline training (Egg, 2021). To do this, GrubHub employed online updates on new samples. Online updates can cause catastrophic forgetting of past knowledge due to concept drift (McCloskey & Cohen, 1989), but due to the rapidly changing preferences of their customers, forgetting past knowledge was desirable. However, for most industry applications, catastrophic forgetting is unacceptable because past knowledge must be retained. For CL to be widely adopted by industry, it must be shown to inhibit model decay in pre-trained models, provide significant computational benefits, work for arbitrary shifts in concepts, and ideally be as effective as offline training. The majority of CL algorithms deviate from these criteria, i.e., they impose constraints on storage that are irrelevant to industry problems, shun pre-trained models, perform worse than offline retraining, design algorithms for only a single concept drift distribution (e.g., class incremental learning), and/or are more computationally expensive than offline retraining (Harun et al., 2023a; b; Prabhu et al., 2023b; a; Verwimp et al., 2024; Lee et al., 2023).
@@
Here, we study CL for mitigating model decay and introducing new concepts into a DNN, where we align the constraints with those of industry where the goal is to deploy effective pre-trained models and update them over time for arbitrary concept shifts. To do this, we initially studied updating ImageNet-1K pre-trained models with new classes in class incremental learning (CIL) experiments. We used computationally constrained cumulative rehearsal to mitigate forgetting, where rehearsal involves mixing old samples with new ones to prevent forgetting, and cumulative rehearsal does this using all previously observed samples. Computational constraints were enforced by limiting the total number of rehearsal updates permitted. However, we found that CL’s stability gap was a major obstacle to our goal (De Lange et al., 2023), as shown in Fig. 1.
@@
The stability gap phenomenon refers to the observation that when the model is updated on new data, accuracy on the classes observed in earlier batches plummets before rehearsal or other CL methods gradually recover old performance. In the typical CIL setting we adopted, samples arrive in batches where each batch contains classes that are only within that batch. As seen in Fig. 1, the performance plummets when learning a new batch during the rehearsal sessions and does not fully recover. In our work, we seek to understand why the stability gap happens in our scenario and how to effectively mitigate it, which would enable using fewer rehearsal updates while achieving higher accuracy.
@@
Figure 1: An overview of stability gap phenomenon. The stability gap is a phenomenon that occurs in CL when learning new data, where accuracy on previously learned data (Y-axis) drops significantly as a function of training iterations when a new distribution is introduced (X-axis). Fig.(a) illustrates this behavior during CIL, where a network pre-trained on ImageNet-1K, learns 365 new classes from Places365-LT over five rehearsal sessions. Each rehearsal session involves 600 training iterations that combine samples from the old and new tasks. A gray dotted vertical line indicates the end of each rehearsal session. When rehearsal begins, accuracy on the old task for the conventional rehearsal drops dramatically before slowly recovering, although it fails to recover the original performance on the old data. Fig.(b) shows the stability gap in the learning curve averaged over five rehearsal sessions. In this work, we seek to understand why the stability gap occurs through a series of hypothesis-driven experiments.
We test two hypotheses for why the stability gap occurs in CIL with pre-trained DNNs:
@@
    The stability gap is caused in part due to having a large loss at the output layer for the new classes. To test this hypothesis, we study two methods to mitigate the large loss  [^1] Referring to the loss when we only train the output layer and freeze all other layers. in the output layer for the new classes. The first method is to initialize the output layer in a data-driven approach rather than randomly initializing the output units responsible for the new classes. The second method is a specialized form of soft targets for the network, rather than the typical hard targets used for the network where these soft targets are designed to improve performance for the new classes while minimally perturbing others.
@@
    The stability gap is caused in part due to excessive network plasticity. We test this hypothesis by controlling the level of plasticity in network layers in a dynamic manner. For hidden layers, we test this hypothesis using LoRA (Hu et al., 2021), which reduces the number of trainable parameters in the hidden layers of the network. For rehearsal methods, after each rehearsal session, these weights are folded into the original network weights. For the output layer, we test this hypothesis by freezing the output units for classes seen in earlier batches during rehearsal.
@@
    We are the first to study overcoming the stability gap. We test the aforementioned hypotheses in CIL and discover that they both play an important role. We propose novel metrics to measure the stability, plasticity, and continual knowledge gaps, which are normalized relative to a jointly trained universal upper bound.
@@
    We are the first to study maintaining or improving performance on ImageNet-1K while learning additional classes, an aspect not investigated in previous research employing ImageNet-1K pre-trained backbones (Wang et al., 2022b; a; Smith et al., 2023; Gao et al., 2023; McDonnell et al., 2024). We study this to learn new data for both the CIL and IID (independent and identically distributed) distributions. These experiments are conducted by combining ImageNet-1K with Places365 or Places365-LT (a combined 1365 total classes).
@@
    We develop a method that greatly mitigates the stability gap, resulting in significant improvements in computational efficiency. Our method requires 16.7×16.7$\times$16.7 × fewer network updates than a jointly-trained (upper bound) model. In terms of TFLOPs, our method provides a 31.9×31.9$\times$31.9 × speedup (see Fig. 5). For the IID CL distribution, the method achieves backward transfer, where learning the new dataset helps to improve ImageNet-1K accuracy.
@@
    We show that our method is effective and can be integrated seamlessly with various CL methods. Specifically, it performs well in conjunction with the rehearsal methods Vanilla Rehearsal, DERpp, GDumb, and REMIND as well as the non-rehearsal method LwF.
@@ 2 Background
A variety of methods have been proposed to learn continuously from non-stationary datasets in continual learning (see Zhou et al. (2023) for review). These methods can be broadly divided into three categories: 1) Rehearsal-based methods store or reconstruct a subset of old data to rehearse alongside new data while learning a new batch (Chaudhry et al., 2019; Hou et al., 2019; Rebuffi et al., 2017; Wu et al., 2019), 2) Regularization-based methods constrain weight updates by adding additional regularization in the loss function (Aljundi et al., 2018; Chaudhry et al., 2018; Dhar et al., 2019; Kirkpatrick et al., 2017), and 3) Parameter-isolation based methods allocate multiple sets of parameters or multiple copies of the model to different incremental batches (Douillard et al., 2021; Yan et al., 2021; Yoon et al., 2020).
While most of the CL research community has focused on mitigating catastrophic forgetting (McCloskey & Cohen, 1989), there is a growing body of literature demonstrating its ability to reduce the amount of compute needed to update the network (Ghunaim et al., 2023; Harun et al., 2023a; b; c; Prabhu et al., 2023b; Verwimp et al., 2024). Recently, a major obstacle to this objective has been identified in CL: The stability gap (De Lange et al., 2023). However, it has been studied on small datasets, e.g., CIFAR using randomly initialized DNNs. This setting is not aligned with studying model decay in industrial settings that involve many-class data streams and preserving performance in large pre-trained DNNs.
@@
Almost all CL methods focus on the catastrophic forgetting problem with evaluations occurring on discrete batch or task transitions (Chaudhry et al., 2018) and fail to capture the stability gap that occurs immediately after a new task is introduced. De Lange et al. (2023) demonstrates the stability gap occurs for a variety of CL methods including rehearsal (Chaudhry et al., 2019), GEM (Lopez-Paz & Ranzato, 2017), EWC (Kirkpatrick et al., 2017)), and LwF (Li & Hoiem, 2017). To quantify the stability gap, they propose continual evaluation metrics based on worst-case performance i.e., the largest drop in accuracy on old batches. However, their metrics do not enable comparing different CL models, because they only quantify the largest drop relative to the same model’s best performance, thus making it impossible to quantify the impact of different training methodologies aimed at mitigating the stability gap. We address this limitation by designing and employing novel metrics that are normalized against a jointly-trained universal upper bound (see Sec. 3.2). We also illustrate this phenomenon using synthetic examples in Appendix F to build intuitions.
We study overcoming the stability gap with both rehearsal and non-rehearsal methods, with a focus on rehearsal due to its effectiveness (van de Ven et al., 2022; Zhou et al., 2023).
@@ 3.1 Formal Continual Learning Setting
To align our work with addressing model decay, we assume CL begins with a pre-trained model capable of $K$-way classification. The goal is to incorporate new data, which may have additional classes, into the model, while preserving or improving performance on the original $K$ classes. For this purpose, we use ImageNet-1K pre-trained models ($K=1000$). While ImageNet-1K pre-trained models are commonly employed in CL systems, previous studies do not attempt to maintain or measure performance on ImageNet-1K itself (Wang et al., 2022b; a; Smith et al., 2023; Gao et al., 2023; McDonnell et al., 2024). Our research uniquely explores this aspect in CL. Upon deployment, the pre-trained model is exposed to a sequence of data batches over $N-1$ learning sessions, i.e., $\{\mathcal{S}_{2},...,\mathcal{S}_{N}\}$, where $\mathcal{S}_{1}$ is the data used for pre-training. The $j$’th session consists of a batch of $m_{j}$ labeled training samples. Each $\mathcal{S}_{j}$ may follow a different class distribution. The total number of samples in the entire sequence is $M=\sum_{j=1}^{N}m_{j}$. When learning new batch $\mathcal{S}_{j}$, the learner can access $\mathcal{S}_{j}$ and any stored data from previous batches $\mathcal{S}_{1:j-1}$. During test time, the learner is evaluated on test data from all seen batches. The batch (or task) identifier $j$ cannot be exploited during test time.
@@ 3.2 Measuring the Stability Gap
$\Omega_{j}$ records CL performance compared to $\mathcal{A}_{joint}$ or $\mathcal{A}_{best}$. After learning all $N$ batches, $\Omega_{j}$ scores are averaged to indicate average performance gain. The first batch is excluded since that is used for pre-training. For all metrics, smaller $\mathcal{S}_{\Delta}$, $\mathcal{P}_{\Delta}$, and $\mathcal{CK}_{\Delta}$ indicate better performance. When $\mathcal{A}_{i}=\mathcal{A}_{joint}$ and $\mathcal{A}_{i}=\mathcal{A}_{best}$ for all $L$ iterations, $\mathcal{S}_{\Delta}$, $\mathcal{P}_{\Delta}$, and $\mathcal{CK}_{\Delta}$ become zero which is desirable. Negative value means knowledge transfer between new and old batches which is also desirable. These metrics are applicable for both offline CL (aka incremental-batch CL) and online CL with any data distributions including CIL and IID.
@@ 4 Stability Gap Mitigation Methods
In this section, we describe the methods we use to test our hypotheses for why the stability gap occurs.
Weight Initialization. In CIL, typically the output units for new classes are randomly initialized causing those units to produce a high loss during backpropagation. We hypothesize that using data-driven initialization for new class units will reduce the loss and therefore reduce the stability gap. To test this, we initialize them to the mean of unit length embeddings for that class, i.e.,
@@
Hard vs. Dynamic Soft Targets. For classification, models are often trained with hard targets, i.e., at training iteration $i$ a one-hot vector $\mathbf{t}_{i}$ with a ‘1’ in position $k$ corresponding to the correct class. We hypothesize hard target training is partially responsible for the stability gap. Intuitively, hard targets are one-hot encoded and enforce strict inter-class independence despite several classes sharing distributional similarities. This property of hard targets, therefore, also causes a large initial loss when learning new classes. Soft targets, on the other hand, can help the network to retain the joint inter-class distributions, which further ameliorates the perturbation of learned classes. To test this, we use soft targets constructed such that the model’s predictions on previously learned classes are largely preserved.
@@
At learning iteration $i$, let $P\left(\mathbf{x}_{i};\theta_{i}\right)$ be the model’s output softmax probabilities for sample $\mathbf{x}_{i}$ from class $k$ given the model’s current parameters $\theta_{i}$ and the predicted class be $y^{\prime}_{i}={\operatorname*{arg\,max}_{k}}P\left({k|{\mathbf{x}}_{i};\theta_{i}}\right)$. We maintain a running average vector $\mathbf{u}_{k}\in\mathbb{R}^{K}$ of the softmax probabilities for each class that is updated when an example from class $k$ is observed, i.e.,
$\mathbf{u}_{k}\leftarrow\frac{c_{k}\mathbf{u}_{k}+P\left(\mathbf{x}_{i};\theta_{i}\right)}{c_{k}+1},$ (5)
@@
where $c_{k}$ is a counter for class $k$ that is subsequently increased by 1, and $\mathbf{u}_{k}$ is initialized to a uniform distribution prior to the running updates. Subsequently, soft targets $\mathbf{t}_{i}$ for iteration $i$ are constructed by setting ${\mathbf{t}}_{i}\leftarrow{\mathbf{u}}_{k}$ and then setting the element for the correct class to 1, i.e., ${\mathbf{t}}_{i}\left[k\right]\leftarrow 1$. If $y^{\prime}_{i}\neq k$, then we also set ${\mathbf{t}}_{i}\left[y^{\prime}_{i}\right]\leftarrow 1/K$. Subsequently, ${\mathbf{t}}_{i}$ is normalized to sum to 1 and used to update the network. This strategy results in targets that minimally perturb the network and smaller loss values.
@@
Limiting Hidden Layer Plasticity Using LoRA. To accumulate knowledge over time, most CL approaches update the entire network. Given that each batch of data in CL is relatively small, we hypothesize that this leads to excessively perturbing hidden representations leading to a larger stability gap. To test this hypothesis, we constrain the number of trainable parameters in hidden representations by using a network adaptor. While there are various network adaptors (Han et al., 2024) to restrict network plasticity, we consider low-rank adaptation (LoRA) (Hu et al., 2021) due to its simplicity and effectiveness. Specifically, we inject LoRA weights into the linear layers of the network, and only these parameters and the output layer are updated greatly reducing the number of trainable parameters.
@@
For batch $j$, let $\mathbf{W}^{j-1}\in\mathbb{R}^{d\times g}$ be a previously learned linear layer. At the start of a each learning session, we reparameterize this layer by replacing $\mathbf{W}^{j-1}$ with
@@
where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times g}$ are the LoRA adapter parameters with rank $r\ll\min(d,g)$. Only $\mathbf{B}$ and $\mathbf{A}$ are plastic, with $\mathbf{A}$ initialized with random Gaussian values and $\mathbf{B}$ initialized to a zero matrix, so $\mathbf{BA}=\mathbf{0}$ at the beginning of the learning session. At the end of the session the LoRA parameters are folded into the network, i.e., $\mathbf{W}^{j}\leftarrow{\mathbf{\Theta}}^{j}$. In LoRA experiments, only the output layer and the LoRA parameters are plastic.
Limiting Output Layer Plasticity via Targeted Freezing. In CIL, large changes in the network’s representations for old classes increase the stability gap. While LoRA restricts plasticity in hidden representations, we hypothesize that restricting plasticity in the output layer could also be helpful for CIL. We therefore explore freezing output layer weights for classes that have been previously learned in earlier batches. For rehearsal methods, samples from classes seen in earlier batches have the hidden layers updated as usual. We refer to this technique as old output class freezing (OOCF).

### v4

@@ Abstract
Pre-trained deep neural networks (DNNs) are being widely deployed by industry for making business decisions and to serve users; however, a major problem is model decay, where the DNN’s predictions become more erroneous over time, resulting in revenue loss or unhappy users. To mitigate model decay, DNNs are retrained from scratch using old and new data. This is computationally expensive, so retraining happens only once performance significantly decreases. Here, we study how continual learning (CL) could potentially overcome model decay in large pre-trained DNNs and greatly reduce computational costs for keeping DNNs up-to-date. We identify the “stability gap” as a major obstacle in our setting. The stability gap refers to a phenomenon where learning new data causes large drops in performance for past tasks before CL mitigation methods eventually compensate for this drop. We test two hypotheses to investigate the factors influencing the stability gap and identify a method that vastly reduces this gap. In large-scale experiments for both easy and hard CL distributions (e.g., class incremental learning), we demonstrate that our method reduces the stability gap and greatly increases computational efficiency. Our work aligns CL with the goals of the production setting, where CL is needed for many applications  [^1] Code is available at https://yousuf907.github.io/sgmsite.
@@ 1 Introduction
Continual learning (CL) is a promising solution for preventing model decay (Huyen, 2022a; b; Jain & Shenoy, 2022), where in CL the goal is to update a DNN incrementally with new data while preserving the previous knowledge (Parisi et al., 2019). In a GrubHub study, this enabled them to avoid model decay and provided a 45$\times$ decrease in training costs compared to daily offline training (Egg, 2021). To do this, GrubHub employed online updates on new samples. Online updates can cause catastrophic forgetting of past knowledge due to concept drift (McCloskey & Cohen, 1989), but due to the rapidly changing preferences of their customers, forgetting past knowledge was desirable. However, catastrophic forgetting is unacceptable for many industry applications because past knowledge must be retained. For CL to be widely adopted by industry, it must be shown to inhibit model decay in pre-trained models, provide significant computational benefits, work for arbitrary shifts in concepts, and ideally be as effective as offline training. The majority of CL algorithms deviate from these criteria, i.e., they impose constraints on storage that are irrelevant to industry problems, shun pre-trained models, perform worse than offline retraining, design algorithms for only a single concept drift distribution (e.g., class incremental learning), and/or are more computationally expensive than offline retraining (Harun et al., 2023a; b; Prabhu et al., 2023b; a; Verwimp et al., 2024; Lee et al., 2023).
@@
Here, we study using CL to mitigate model decay and to introduce new concepts into a DNN. Initially, we studied updating ImageNet-1K pre-trained models with new classes in class incremental learning (CIL) experiments. We used computationally constrained cumulative rehearsal to mitigate forgetting, where rehearsal involves mixing old samples with new ones to prevent forgetting, and cumulative rehearsal does this using all previously observed samples. Computational constraints were enforced by limiting the total number of rehearsal updates permitted. However, we found that CL’s stability gap was a major obstacle to our goal (De Lange et al., 2023), as shown in Fig. 1.
@@
The stability gap refers to the observation that when a model is updated on new data, accuracy on the classes observed in earlier batches plummets before rehearsal or other CL methods gradually recover old performance. In the typical CIL setting we adopted, samples arrive in batches where each batch contains classes only within that batch. As seen in Fig. 1, the performance plummets when learning a new batch during the rehearsal sessions and does not fully recover. The stability gap can be seen as transient forgetting due to introducing a new task. Typically, catastrophic forgetting refers to the performance drops observed at the end of learning new tasks (at task transitions). In contrast, the stability gap refers to the performance drops observed over learning steps (between task transitions). In our work, we seek to understand why the stability gap happens in our scenario and how to effectively mitigate it, enabling using fewer rehearsal updates while achieving higher accuracy.
@@
Figure 1: An overview of stability gap phenomenon. The stability gap is a phenomenon that occurs in CL when learning new data, where accuracy on previously learned data (Y-axis) drops significantly as a function of training iterations when a new distribution is introduced (X-axis). Fig.(a) illustrates this behavior during CIL, where a network pre-trained on ImageNet-1K, learns 365 new classes from Places365-LT over five rehearsal sessions. Each rehearsal session involves 600 iterations that combine samples from the old and new tasks. A gray dotted vertical line marks the end of a rehearsal session or a task transition. When rehearsal begins, accuracy on the old task for the conventional rehearsal drops dramatically before slowly recovering, although it fails to recover the original performance on the old data. The traditional measures of catastrophic forgetting focus on performance at task transitions (red diamonds), ignoring significant forgetting that occurs during the learning process between task transitions. Fig.(b) shows the stability gap in the learning curve averaged over five rehearsal sessions. In this work, we attempt to mitigate the stability gap.
We test two hypotheses to examine the stability gap in CIL with pre-trained DNNs:
@@
    The stability gap is increased in part due to having a large loss at the output layer for the new classes. To test this hypothesis, we study two methods to mitigate the large loss  [^2] Referring to the loss when we only train the output layer (i.e., final classification layer) and freeze all other layers. in the output layer for the new classes. The first method is to initialize the output layer in a data-driven approach rather than randomly initializing the output units responsible for the new classes. The second method is a specialized form of soft targets for the network, rather than the typical hard targets used for the network, where these soft targets are designed to improve performance for the new classes while minimally perturbing others.
@@
    The stability gap is increased in part due to excessive network plasticity. We test this hypothesis by controlling the level of plasticity in network layers in a dynamic manner. For hidden layers, we test this hypothesis using LoRA (Hu et al., 2021), which reduces the number of trainable parameters in the hidden layers of the network. For rehearsal methods, after each rehearsal session, these weights are folded into the original network weights. For the output layer, we test this hypothesis by freezing the output units for classes seen in earlier batches during rehearsal.
@@
    We are the first to study overcoming the stability gap. We test the aforementioned hypotheses in CIL and discover that they both play an important role. We propose novel metrics to measure the stability, plasticity, and continual knowledge gaps.
@@
    We are the first to study maintaining or improving performance on ImageNet-1K while learning additional classes, an aspect not investigated in previous research employing ImageNet-1K pre-trained backbones (Wang et al., 2022b; a; Smith et al., 2023; Gao et al., 2023; McDonnell et al., 2024). We study this for both the CIL and IID (independent and identically distributed) distributions. These experiments are conducted by combining ImageNet-1K with Places365 or Places365-LT (a combined 1365 total classes).
@@
    We develop a method that greatly mitigates the stability gap and significantly improves computational efficiency. Our method requires 16.7×16.7$\times$16.7 × fewer network updates than a jointly-trained (upper bound) model. In terms of TFLOPs, our method provides a 31.9×31.9$\times$31.9 × speedup (see Fig. 5). For the IID CL distribution, the method achieves backward transfer, where learning the new dataset helps to improve ImageNet-1K accuracy.
@@
    We show that our method is effective and we integrate it into existing CL methods. Specifically, it performs well in conjunction with the rehearsal methods Vanilla Rehearsal, DERpp, GDumb, and REMIND as well as the non-rehearsal method LwF.
@@ 2 Background
Many methods have been proposed to learn continuously from non-stationary datasets in continual learning (see Zhou et al. (2023) for review). These methods can be broadly divided into three categories: 1) Rehearsal-based methods store or reconstruct a subset of old data to rehearse alongside new data while learning a new batch (Chaudhry et al., 2019; Hou et al., 2019; Rebuffi et al., 2017; Wu et al., 2019), 2) Regularization-based methods constrain weight updates by adding additional regularization in the loss function (Aljundi et al., 2018; Chaudhry et al., 2018; Dhar et al., 2019; Kirkpatrick et al., 2017), and 3) Parameter-isolation based methods allocate multiple sets of parameters or multiple copies of the model to different incremental batches (Douillard et al., 2021; Yan et al., 2021; Yoon et al., 2020). Since parameter isolation methods do not allow backward transfer, the stability gap does not apply to them De Lange et al. (2023).
While most of the CL research community has focused on mitigating catastrophic forgetting (McCloskey & Cohen, 1989), there is a growing body of literature demonstrating its ability to reduce the amount of compute needed to update the network (Ghunaim et al., 2023; Harun et al., 2023a; b; c; Prabhu et al., 2023b; Verwimp et al., 2024). Recently, a major obstacle to this objective has been identified in CL: The stability gap (De Lange et al., 2023). However, it has been studied on small datasets, e.g., CIFAR, using randomly initialized DNNs. This setting is not aligned with studying model decay in industrial settings that involve many-class data streams and preserving performance in large pre-trained DNNs.
@@
Almost all CL methods focus on the catastrophic forgetting problem with evaluations occurring on discrete batch or task transitions (Hayes et al., 2018; Chaudhry et al., 2018) and fail to capture the stability gap that occurs immediately after a new task is introduced (Fig. 1). De Lange et al. (2023) demonstrates the stability gap occurs for a variety of CL methods, including rehearsal (Chaudhry et al., 2019), GEM (Lopez-Paz & Ranzato, 2017), EWC (Kirkpatrick et al., 2017)), and LwF (Li & Hoiem, 2017). To quantify the stability gap, they propose continual evaluation metrics based on worst-case performance i.e., the largest drop in accuracy on old batches. However, their metrics do not enable comparing different CL models because they only quantify the largest drop relative to the same model’s best performance, thus making it impossible to quantify the impact of different training methodologies aimed at mitigating the stability gap. We address this limitation using novel metrics that are normalized against a jointly-trained universal upper bound (see Sec. 3.2). We also illustrate this phenomenon using synthetic examples in Appendix F to build intuitions. We study overcoming the stability gap with both rehearsal and non-rehearsal methods, with a focus on rehearsal due to its effectiveness (van de Ven et al., 2022; Zhou et al., 2023).
@@ 3.1 Formal Continual Learning Setting
To align our work with addressing model decay, we assume CL begins with a pre-trained model capable of $K$-way classification. The goal is to incorporate new data, which may have additional classes, into the model, while preserving or improving performance on the original $K$ classes. For this purpose, we use ImageNet-1K pre-trained models ($K=1000$). While ImageNet-1K pre-trained models are commonly employed in CL systems, previous studies do not attempt to maintain or measure performance on ImageNet-1K dataset itself (Wang et al., 2022b; a; Smith et al., 2023; Gao et al., 2023; McDonnell et al., 2024). Our research uniquely explores this aspect in CL. Upon deployment, the pre-trained model is exposed to a sequence of data batches over $N-1$ learning sessions, i.e., $\{\mathcal{S}_{2},\cdots,\mathcal{S}_{N}\}$, where $\mathcal{S}_{1}$ is the data used for pre-training. The $j$’th session consists of a batch of $m_{j}$ labeled training samples, i.e., $\mathcal{S}_{j}=\{(\mathbf{x}_{j},y_{j})_{i}\}_{i=1}^{m_{j}}$, where $\mathbf{x}_{j}$ is an instance of category $y_{j}\in Y_{j}$ and $Y_{j}$ is the label space of task $j$. In the CIL setting, each $\mathcal{S}_{j}$ contains non-overlapping classes, i.e., $Y_{j}\cap Y_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$. The total number of samples in the entire sequence is $M=\sum_{j=1}^{N}m_{j}$. When learning new batch $\mathcal{S}_{j}$, the learner can access $\mathcal{S}_{j}$ and any stored data from previous batches $\mathcal{S}_{1:j-1}$. During test time, the learner is evaluated on test data from all seen classes, $\mathcal{Y}_{j}=Y_{1}\cup\cdots Y_{j}$. The batch (or task) identifier $j$ cannot be exploited during test time.
@@ 3.2 Measuring the Stability Gap
$\Omega_{j}$ records CL performance compared to $\mathcal{A}_{joint}$ or $\mathcal{A}_{best}$. After learning all $N$ batches, $\Omega_{j}$ scores are averaged to indicate average performance gain. The first batch is excluded since that is used for pre-training. For all metrics, smaller $\mathcal{S}_{\Delta}$, $\mathcal{P}_{\Delta}$, and $\mathcal{CK}_{\Delta}$ indicate better performance. When $\mathcal{A}_{i}=\mathcal{A}_{joint}$ and $\mathcal{A}_{i}=\mathcal{A}_{best}$ for all $L$ iterations, $\mathcal{S}_{\Delta}$, $\mathcal{P}_{\Delta}$, and $\mathcal{CK}_{\Delta}$ become zero which is desirable. A negative value means knowledge transfer between new and old batches, which is desirable. These metrics apply for offline CL (i.e., incremental-batch CL) and online CL with any data distributions, including CIL and IID.
@@ 4 Stability Gap Mitigation Methods
This section describes the methods we use to test our hypotheses regarding the factors that increase the stability gap. We use our evaluation metrics, $\mathcal{S}_{\Delta}$, $\mathcal{P}_{\Delta}$, and $\mathcal{CK}_{\Delta}$ to validate the efficacy of proposed mitigation methods. We include additional discussion in Appendix H.
Weight Initialization. In CIL, the output units for new classes are typically randomly initialized causing those units to produce a high loss during backpropagation. We hypothesize that using data-driven initialization for new class units will reduce the loss and therefore reduce the stability gap, $\mathcal{S}_{\Delta}$. To test this, we initialize them to the mean of unit length embeddings for that class, i.e.,
@@
Hard vs. Dynamic Soft Targets. For classification, models are often trained with hard targets, i.e., at training iteration $i$ a one-hot vector $\mathbf{t}_{i}$ with a ‘1’ in position $k$ corresponding to the correct class. We hypothesize hard target training is partially responsible for the stability gap. Intuitively, hard targets are one-hot encoded and enforce strict inter-class independence despite several classes sharing distributional similarities. This property of hard targets, therefore, also causes a large initial loss when learning new classes. Soft targets, on the other hand, can help the network retain the joint inter-class distributions, which further ameliorates the perturbation of learned classes. To test this, we use soft targets constructed such that the model’s predictions on previously learned classes are largely preserved.
@@
At learning iteration $i$, let $P\left({k|{\mathbf{x}}_{i};\theta_{i}}\right)$ be the model’s output softmax probabilities for sample $\mathbf{x}_{i}$ from class $k$ given the model’s current parameters $\theta_{i}$ and the predicted class be $y^{\prime}_{i}={\operatorname*{arg\,max}_{k}}P\left({k|{\mathbf{x}}_{i};\theta_{i}}\right)$. We maintain a running average vector $\mathbf{u}_{k}\in\mathbb{R}^{K}$ of the softmax probabilities for each class that is updated when an example from class $k$ is observed, i.e.,
$\mathbf{u}_{k}\leftarrow\frac{c_{k}\mathbf{u}_{k}+P\left({k|{\mathbf{x}}_{i};\theta_{i}}\right)}{c_{k}+1},$ (5)
@@
where $c_{k}$ is a counter for class $k$ that is subsequently increased by 1, and $\mathbf{u}_{k}$ is initialized to a uniform distribution prior to the running updates. Subsequently, soft targets $\mathbf{t}_{i}$ for iteration $i$ are constructed by setting ${\mathbf{t}}_{i}\leftarrow{\mathbf{u}}_{k}$ and then setting the element for the correct class to 1, i.e., ${\mathbf{t}}_{i}\left[k\right]\leftarrow 1$. If $y^{\prime}_{i}\neq k$, then we also set ${\mathbf{t}}_{i}\left[y^{\prime}_{i}\right]\leftarrow 1/K$. Subsequently, ${\mathbf{t}}_{i}$ is normalized to sum to 1 and used to update the network. This strategy results in targets that minimally perturb the network and smaller loss values. In Appendix. G, we illustrate the process of updating soft targets.
@@
Limiting Hidden Layer Plasticity Using LoRA. To accumulate knowledge over time, most CL approaches update the entire network. Given that each batch of data in CL is relatively small, we hypothesize that this leads to excessively perturbing hidden representations, leading to a larger stability gap. To test this hypothesis, we constrain the number of trainable parameters in hidden representations using a network adaptor. While there are various network adaptors (Han et al., 2024) that restrict network plasticity, we use low-rank adaptation (LoRA) (Hu et al., 2021) due to its simplicity and effectiveness. Specifically, we inject LoRA weights into the linear layers of the network, and only these parameters and the output layer are updated, which greatly reduces the number of trainable parameters.
@@
For batch $j$, let $\mathbf{W}^{j-1}\in\mathbb{R}^{d\times g}$ be a previously learned linear layer. At the start of each learning session, we reparameterize this layer by replacing $\mathbf{W}^{j-1}$ with
@@
where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times g}$ are the LoRA adapter parameters with rank $r\ll\min(d,g)$. Only $\mathbf{B}$ and $\mathbf{A}$ are plastic, with $\mathbf{A}$ initialized with random Gaussian values and $\mathbf{B}$ initialized to a zero matrix, so $\mathbf{BA}=\mathbf{0}$ at the beginning of the learning session. At the end of the session, the LoRA parameters are folded into the network, i.e., $\mathbf{W}^{j}\leftarrow{\mathbf{\Theta}}^{j}$. In LoRA experiments, only the output layer and the LoRA parameters are plastic.
Limiting Output Layer Plasticity via Targeted Freezing. In CIL, large changes in the network’s representations for old classes increase the stability gap. While LoRA restricts plasticity in hidden representations, we hypothesize that restricting plasticity in the output layer could also be helpful for CIL. Therefore, we study freezing output layer weights for classes previously learned in earlier batches. For rehearsal methods, samples from classes seen in earlier batches have the hidden layers updated as usual. We refer to this technique as old output class freezing (OOCF).